OpenStudy (anonymous):

What adds up to "9" and multiply to "-36"?

OpenStudy (jdoe0001):

\(\bf 36\implies 2\cdot 2\cdot 3\cdot 3\) what do you think?

OpenStudy (jdoe0001):

\(\bf 36\implies 2\cdot 2\cdot 3\cdot 3\\ \quad \\ (2\cdot 2\cdot 3)-(3)\implies \square ?\)

OpenStudy (anonymous):

12 and -3 thank you @jdoe0001

OpenStudy (jdoe0001):

yw